# Keys to Robust Edits: From Theoretical Insights to Practical Advances

### v1

@@ 5.1 Experimental Setup
Following previous work, we use the CounterFact (Meng et al., 2023a) datasets, choosing LLaMA-2 as our base model. In addition to the prompt from CounterFact dataset, we additionally consider three types of perturbation in our experiments, namely prompt appended with unrelated long context, subject rephrase and random shuffled subject. Even though the shuffled subject does not contain the same semantic meaning, it demonstrates how keys shift when the position of same word occurs at different positions.
We collect 10 rephrases for each subject by prompting gpt-4o-mini. The prompt we use can be found in Appendix. For long context, we follow Ma et al. (2024c) and extract random text span of 300 tokens from Wikitext-103 Merity et al. (2016). For rephrased prompts, we use the paraphrases of prompts released by Patil et al. (2023). For shuffled subject, we sample 10 random orderings of tokens in the subject. We use 100 samples in our valid set for empirical analyses.
@@ 5.2 Empirical Statistics of Keys, Values and Others
Dissimilar Keys
In Figure 3 (left), we present the distribution of whitening dot products $k_{1}C^{-1}k_{2},\forall k_{1},k_{2}\in K_{s}$. For each subject in CounterFact, we compute the dot product for each pair of keys of a subject’s rephrases. We utilize the inputs to the FFN’s down projection of layer 5 of LLaMA-2 as our keys, consistent with previous ROME experiments. Additionally, we include the dot product values of randomly sampled keys as a baseline for comparative analysis. Higher values indicate greater similarity between key pairs in the whitening space.
Analysis of the distribution reveals notable findings – While some key pairs exhibit high similarity, a substantial portion of the distribution mass concentrates near zero, aligning with the random keys baseline. This suggests a significant variability in the similarity of rephrased subject representations. Intriguingly, some key pairs demonstrate negative similarity. This phenomenon warrants further investigation to understand its implications for semantic representation and model behavior.
These findings challenges the intuition that semantically equivalent subjects should have similar representations, and poses severe challenges to the effectiveness of edits. When the edited key has near-zero or negative similarity with other keys, based on Lemma 3 it becomes virtually impossible for the edited value to be retrieved, potentially compromising the robustness of the edit.
@@
Figure 4: Left: CounterFact subjects have unrelated prefixs which are close in keys. Right: Semantically similar subjects brings challenges to specificity.
@@
We also investigate whether there exists different subjects that have highly similar keys. To this end, we iterate through a slice of Wikitext-103 dataset (8̃0M tokens) and select those are close to subjects in CounterFact in the whitening space. We filter those tokens whose prefix has the same subject token and collect the top-10 unrelated keys of each subject in CounterFact. The left of Figure 4 plots the distribution of whitening similarities between unrelated prefixs and CounterFact subjects. We find that a large portion of them has extremely high whitening similarity score, i.e., > 2500. Based on our theory, it indicates that any edit affects these subjects would inevitably affect the output on these unrelated prefixs.
On the right side of Figure 4, we show cases of subjects and their top-1 prefix in terms of whitening similarities. Interestingly, we observe that a subject can exhibit similarity in distributional semantics (Lenci & Sahlgren, 2023) to its corresponding top unrelated prefix. For example, the keys of Michael Jordan are highly similar to keys of a prefix related to Kobe. Considering that these two basketball players has much in common in many perspectives, it makes sense that their keys are similar. However, an edit to Michael Jordan affects Kobe would be definitely unreasonable.
Loud Voices.
Figure 5: Values before and after edit with ROME.
In Figure 5, we present the distribution of values before and after edits, using LLaMA-2 7B and ROME. The results demonstrate that post-edit values exhibit significantly larger L2 norms compared to pre-edit values. This observation aligns with our findings in Lemma 4.4 and 3, which suggest that edited values must be sufficiently large to effect changes on the current key and influence distant keys.
Summary.
@@ 6 Robust Edit Pathway
Figure 6: Architecture design of the adaptor.
Our idea to address above issues is to disentangle the keys from the models’ inner representations, by introducing a potential branching path as keys of edited facts.
We achieve this by placing an adaptor directly after the keys and modify the representations of the keys when it is necessary. As shown in Figure 6, our adaptor consists of two modules, a projection module that is responsible for aligning the keys and a gate module that activates the adaptor when a token representation needs to be edit:
@@
We train the adaptor by aggregating the keys of same subject $k_{s}\in K_{s}$ toward our injected target key $k_{*}$:
$\mathcal{L}=-{|(\frac{\hat{k_{s}}}{||\hat{k_{s}}||_{2}})^{T}C^{-1}k_{*}|}$ (6)
where $\hat{k_{s}}$ is the output keys after adaptor. The intuition is inspired by Lemma 3 and 4.7. If the edited key is close to the keys of the same subejct, especially those we found dissimilar in Section 5.2, the edit would be more robust.
In practice, we find that model inclines to ‘cheat’ by simply increasin the norm of $k_{s}$ and thus we normalize the output of $f$. In practice, we take the last token of rephrased subjects over different contexts and rephrased templates as $k_{s}$. This objective, built on the whitening similarity, further strengthen the validness of our theorectical results.
For testing, we use a gate threshold $\tau$ to determine whether to activate this branch. This gate mechanism allows the model to dynamically decide whether the original keys should be modified. If not, the keys are left intact and thus ensure the locality of edits.
Algorithm 1 illustrates the training and testing with our adaptor.
@@ 6.1 Experimental Results
Table 1: REP improves over both in-domain and out-of-domain robustness metrics with $\tau=0.9$.
Edit Performance In-Domain Out-of-Domain
Success Locality Rephrase Shuffle Long Rephrase Shuffle Long
LLaMA2-7B ROME 84.5 96.5 52.1 11.4 74.6 55.7 5.9 74.5
ROME + REP 81.0 91.7 79.1 75.3 78.6 71.8 43.6 79.8
Mistral-7B ROME 94.5 94.2 65.6 14.2 86.7 65.8 14.5 86.0
ROME + REP 91.5 91.1 90.9 86.0 88.4 83.9 48.8 86.8
@@
We evaluate our Robust Edit Pathway with a representative locate-and-edit method, namely ROME. We use the LLaMA2-7B model as our base model and CounterFact as our dataset. We filter knowledge triplets of CounterFact not presented in the model as Meng et al. (2023a) did, and randomly sample 100 knowledge triplets as the validation set and 400 triplets as the test set. Our research concentrates on the single-edit paradigm, wherein one fact is injected into the LLM at a time. While other studies in model editing explore modifying multiple facts continuously (Mitchell et al., 2022c; Hartvigsen et al., 2023; Meng et al., 2023b), we have found that robustly injecting even a single fact presents significant challenges. Therefore, we maintain our focus on this fundamental aspect in this work.
For evaluation, we consider the following metrics: (1) Success: the ratio of targeted knowledge achieving the top probability; (2) Locality: the raito of related but non-identical facts kept intact by the edit; (3) Rephrased: the ratio of same subject with rephrased surface form achieving a success edit; (4) Shuffled Subject: the ratio of shuffled form of subjects achieving a success edit; (5) Long Context: the ratio of facts with random context appended before achieving a success edit.
Metrics (1) and (2) are widely used in previous studies (Meng et al., 2023a; b) of knowledge editing methods, where the success rate indicates the effectiveness and the locality represents the specificity of editing methods. Metrics (3)-(5) are robustness metrics aligned with our previous analyses. Improving these metrics suggests a more robust editing method. We report robustness metrics at both in-domain, where the test cases are seen in training adaptor, and out-of-domain, where the test cases are not seen by adaptor. Note that in our ‘in-domain’, we do not reveal the target knowledge to the model, we only aggregate the keys.
@@
Our resutls are shown in Table 1. Our baseline methods, ROME, achieve solid edit success rate (8̃5%) and good locality scores (>90%) on both LLaMA2-7B and Mistral-7B. Nonetheless, these methods are easy to fail with robustness tests. ROME’s success rate drops 3̃0% with rephrased subjects, drops 7̃0% with shuffled subjects ordering and drops 1̃0% with randomly appended long context. These results reconcile with those reported in previous studies (Ma et al., 2024c).
@@
Then, with a slight cost of Success and Locality, the robustness metrics siginificantly improve when adopting our method. Specifically, after aggregating the keys, the success rate of in-domain tests all reaches about 75-79%, which is close to the success rate without any perturbation (the upper bound). This further validates our theoretical and empirical results. Then, our method also improves over out-of-domain robustness tests, indicating that our adaptor learns to generalize across types of perturbations.
Study of Gate Threshold.
Gate threshold $\tau$ is crucial to the performance of REP. Figure 7 varies different values of $\tau$ and evaluate our metrics. We find that a larger $\tau$ leads to a better locality and success rate. Meanwhile, the robustness metrics first plateau then degrade with the increase of $\tau$, indicating a trade-off between robustness and edit performance.
Figure 7: Validation performance of edited LLaMA2-7B against $\tau$ is plotted.
7 Conclusion and Future Directions
In this work, we have presented a comprehensive analysis of the challenges faced by knowledge editing techniques in large language models, specifically focusing on the robustness issues in locate-and-edit methods. Our theoretical framework, centered on the ’group discussion’ model, provides valuable insights into the mechanisms of knowledge insertion and the requirements for robust and specific edits.
The proposed Robust Edit Pathway (REP) addresses the key limitations identified in our empirical analysis, effectively disentangling editing keys from the LLM’s internal representations. Our extensive evaluations on LLaMA2-7B and Mistral-7B demonstrate significant improvements in robustness across various metrics, including rephrased subjects, shuffled token ordering, and long context additions, while maintaining high success rates and locality.
These results not only validate our theoretical foundations but also offer a promising direction for enhancing the reliability and flexibility of knowledge editing in LLMs. The REP’s ability to generalize across different types of perturbations represents a substantial step forward in addressing the critical robustness issues prevalent in existing techniques.
Future work could explore the scalability of our approach to larger models and more diverse datasets, as well as investigate potential applications in continual edits and dynamic knowledge updating for LLMs. Overall, this research contributes to the ongoing efforts to create more adaptable and trustworthy language models, paving the way for their responsible deployment in real-world applications.
Ethic Statement
We honor the ICLR Code of Ethics. No private data or non-public information was used in this work.
Reproduction Statement
We have appended the data and code in supplementary files for review and reproduction.
@@ Appendix A Proofs
A.3 Proof of Lemma 3 and 4.7
Based on our result above, we can further derive results for
@@ B.1 Data Construction
The unrelevant long contexts are extracted from the Wikitext-103 dataset (Merity et al., 2016). The shuffled tokens are generated via sampling different word ordering. Finally, we filter the samples that are not present in the current LLM, that is, given the prefix, the target tokens are not predicted by the LLMs with the top-1 probabilities. We sample 100 samples for validation and 400 samples for test. To evaluate in-domain and out-of-domain robustness, we split the all three kinds of robustness queries in a 50-50 manner. For each sample, we have 5 in-domain queries and 5 out-of-domain queries.
B.2 Details of Training BPA

### v2

@@ 5.1 Experimental Setup
Following previous work, we use the CounterFact (Meng et al., 2022a) datasets, choosing LLaMA-2 as our base model. In addition to the prompt from CounterFact dataset, we additionally consider three types of perturbation in our experiments, namely prompt appended with unrelated long context, subject rephrase and random shuffled subject. Even though the shuffled subject does not contain the same semantic meaning, it demonstrates how keys shift when the position of same token occurs at different positions.
We collect 10 rephrases for each subject by prompting gpt-4o-mini. The prompt we use can be found in Appendix. For long context, we follow Ma et al. (2024c) and extract random text span of 512 tokens from Wikitext-103 Merity et al. (2016). For rephrased prompts, we use the paraphrases of prompts released by Patil et al. (2023). For shuffled subject, we sample 10 random orderings of tokens in the subject. We use 100 samples in our valid set for empirical analyses.
@@ 5.2 Empirical Statistics of Keys, Values and Others
Dissimilar Keys.
In Figure 3, we present the distribution of whiten similarity $\beta$ for three operations over the original edit along with a random key baseline. The implementation detail can be found in Appendix B.2.
We can see that the similarity after these operations drops drastically. Rephrasing and shuffling word orders generally reduce the similarity from 1.0 to less than 0.4, even to the random level. Appending long context is less destructive, but still reduces the key similarity to [0.2, 0.9]. These results indicate a violation of the robustness requirement, showing a significant variability in the representation of the same subject, making locate-and-edit difficult to retrieve the edited value to be retrieved.
These findings challenges the intuition that semantically equivalent subjects should have similar representations, and poses severe challenges to the effectiveness of edits.
@@
Figure 4: Left: CounterFact subjects have unrelated prefixes which are close in keys. The red dashed line indicates random keys baseline. Right: Semantically similar subjects bring challenges to specificity.
@@
We also investigate whether there exist different subjects that have highly similar keys. To this end, we iterate through a slice of Wikitext-103 dataset (about 80M tokens) and select those close to subjects in CounterFact in the whitening space. We filter those tokens whose prefix has the same subject token and collect the top 10 unrelated keys of each subject in CounterFact. The left of Figure 4 plots the distribution of whitening similarities between unrelated prefixes and CounterFact subjects. We find that a large portion of them has an extremely high whitening similarity score, i.e., > 2500. Based on our theory, it indicates that any edit that affects these subjects would inevitably affect the output on these unrelated prefixes.
On the right side of Figure 4, we present a list of subjects and their top-1 prefix in terms of whitening similarities. Interestingly, we observe that a subject can exhibit similarity in distributional semantics (Lenci and Sahlgren, 2023) to its corresponding top unrelated prefix. For example, the keys of Michael Jordan are highly similar to keys of a prefix related to Kobe. Considering that these two basketball players has much in common in many perspectives, it makes sense that their keys are similar. However, an edit to Michael Jordan affects Kobe would be definitely unreasonable.
@@ 6 Robust Edit Pathway
Our solution is to separate keys from the model’s internal representations by introducing a potential branching path as keys for edited facts.
This is done by adding an adapter after the keys, allowing their representations to be modified when needed. As shown on the right of Figure 2, our adapter consists of two modules, a projection module that is responsible for aligning the keys and a gate module that activates the adapter when a token representation needs to be edited:
@@
We train the adapter by aggregating the keys of same subject $k_{s}\in\mathcal{K}_{s}$ toward our injected target key $k_{*}$:
$\mathcal{L}_{\text{agg}}=-{|(\frac{\hat{k_{s}}}{||\hat{k_{s}}||_{2}})^{T}C^{-1}k_{*}|}$ (4)
where $\hat{k_{s}}$ is the output keys after adapter. The intuition is inspired by Lemma 4.6 and 4.7. If the edited key is close to the keys of the same subject, especially those we found dissimilar in Section 5.2, the edit would be more robust.
In practice, we find that the model inclines to ‘cheat’ by simply increasing the norm of $k_{s}$, and thus we normalize the output of $f$. In practice, we take the last token of rephrased subjects over different contexts and rephrased templates as $k_{s}$. This objective, built on the whiten similarity, further strengthens the validness of our theoretical results.
To address the drift of the target key $k_{*}$ during optimization, we introduce a target consistency loss:
$\mathcal{L}_{\text{consistency}}=MSE(\hat{k_{*}},k_{*})$ (5)
The final training objective combines both components:
$\mathcal{L}=\mathcal{L}_{\text{agg}}+\alpha\mathcal{L}_{\text{consistency}}$ (6)
with $\alpha$ controlling the trade-off.
For testing, we use a gate threshold $\tau$ to determine whether to activate this projection. This gate mechanism allows the model to dynamically decide whether the original keys should be modified. If not, the keys are left intact and thus ensure the locality of edits. The whole algorithm can be found in Appendix.
Method Edit Performance In-Domain Out-of-Domain
[…]
Baseline Methods
ROME 100.0 96.1 63.8 587.4 61.0 13.0 89.8 62.6 13.7 89.8
MEMIT 99.3 91.2 71.9 571.4 73.3 30.0 92.3 73.4 32.0 94.3
R-ROME 99.7 95.8 62.1 583.8 58.9 14.7 89.5 61.7 16.1 90.7
EMMET 99.7 93.8 63.0 584.0 59.7 16.3 83.7 60.9 16.5 83.0
With REP
ROME 100.0+0.0 94.6-1.5 66.9+3.1 587.5+0.1 88.0+27.0 59.9+46.9 91.7+1.9 75.5+12.9 28.7+15.0 91.3+1.5
MEMIT 99.4+0.1 90.8-0.4 74.2+2.3 567.2-4.2 89.9+16.6 58.9+28.9 93.6+1.3 84.4+11.0 45.2+31.5 94.2-0.1
R-ROME 99.9+0.2 94.7-1.1 67.4+5.3 586.0+2.2 88.8+29.9 60.3+45.6 92.0+2.5 76.5+14.8 29.5+13.4 92.0+1.3
EMMET 99.8+0.1 92.2-1.6 68.4+5.4 584.6+0.6 94.4+34.7 82.7+66.4 88.4+4.7 82.9+22.0 42.5+26.0 88.6+5.6
Table 1: The main results of REP across three seeds comparing ROME, MEMIT, R-ROME, and EMMET editing methods on Llama2-7B on CounterFact dataset. REP consistently enhances model performance Results averaged over three seeds with $\tau=0.9$. The upperscript numbers denote the improvement after using REP.
@@ 6.1 Experimental Results
We evaluate our Robust Edit Pathway with representative locate-and-edit methods, namely ROME, MEMIT, R-ROME, and EMMET. We use the LLaMA2-7B, Mistral-7B, and Qwen2-7B as our base model and CounterFact and ZsRE as our datasets. We filter knowledge triplets of datasets not presented in the model as Meng et al. (2022a) did, and randomly sample 100 knowledge triplets as the validation set and 400 triplets as the test set. While other studies in model editing explore modifying multiple facts continuously (Mitchell et al., 2022; Hartvigsen et al., 2024; Meng et al., 2022b), we have found that robustly injecting even a single fact presents significant challenges. Therefore, we keep our focus on single-edit paradigm.
For evaluation, we first follow (Meng et al., 2022a, b) and utilize the following four metrics for edit performance: (1) Success: the ratio of targeted knowledge achieving the top probability; (2) Locality: the ratio of related but non-identical facts kept intact by the edit; (3) Paraphrase (Para.): the ratio of targeted knowledge achieving success on paraphrased prompts; (4) Fluency: the weighted average of bi- and tri-gram entropies.
Moreover, we report the success rate for three robustness tests: paraphrasing subjects, shuffling subjects’ token ordering, and appending long context, as discussed throughout the paper. Improving these metrics suggests a more robust editing method. We report robustness metrics at both in-domain, where the test cases are seen in training adapter, and out-of-domain, where the test cases are not seen by adapter. Note that in our ‘in-domain’, we do not reveal the target knowledge to the model, we only aggregate the keys.
@@
Our results are shown in Table 1. We can see that our baseline methods, ROME, MEMIT, R-ROME, and EMMET achieve near-perfect edit success rates (>99%) while preserving good locality scores (93-96%). Nonetheless, these methods are prune to robustness tests. Taking ROME as an example, the success rate drops 39% with rephrased subjects, 87% with shuffled subjects ordering, and 10.2% with randomly appended long context. These results reconcile with those reported in previous studies (Ma et al., 2024c).
@@
REP improves the robustness of each of the locate-and-edit methods significantly, with a slight cost of locality drop. For instance, REP improves ROME over three robustness tests with +27.0%/+46.9%/+1.9% for in-domain queries, and +12.9%/+15.0%/+1.5% for out-of-domain queries. We also conduct experiments over a different dataset (ZsRE) and two additional base models (QWen and Mistral). The results are shown in Appendix and consistently demonstrate the effectiveness of REP. This validates our theoretical results and empirical insights.
Ablation Studies.
Gate threshold $\tau$ and consistency loss weight $\alpha$ are crucial to the performance of REP. We study them in the Appendix with Figure 5 and Figure 6. We find that a larger $\tau$ and a larger $\alpha$ leads to a better locality and success rate. Meanwhile, the robustness metrics first plateau, then degrade with the increase of $\tau$ and $\alpha$, indicating a trade-off between robustness and edit performance. Throughout our experiments, we use $\tau=0.9$ and $\alpha=5e+4$.
7 Conclusion
In this work, we challenge a core assumption in the locate-and-edit mechanism – the model’s inner representations can serve as semantic keys for editing. We present theoretical results and empirical analyses revealing that these keys are both sensitive and unspecific. To address this issue, we propose the Robust Edit Pathway (REP), which disentangles the editing keys from native model representations. By extensive experiments, we show that REP can significantly enhance robustness over various locate-and-edit methods while maintains the edit success rate.
8 Limitations
While REP demonstrates significant improvements in knowledge editing robustness, our work is limited in the following aspects: (1) REP requires additional training steps to learn the adapter parameters, introducing computational overhead compared to direct editing methods. (2) Our current evaluation focuses on single-fact editing. The effectiveness of REP in scenarios involving multiple interrelated facts or continuous editing remains to be investigated. (3) In this work, we focus on locate-and-edit methods. Even though it is the dominant line of methods in model editing methods, there are still other model editing methods and REP does not apply to them.
9 Acknowledgement
This publication has emanated from research conducted with the financial support of the National Natural Science Foundation of China Key Program under Grant Number 6233000066
@@ Appendix A Proofs
A.3 Proof of Lemma 4.6
For an edit to be robust, it must propagate correctly to all semantically equivalent inputs. We derive this requirement step by step:
1) From Lemma 4.4, a successful edit requires:
$(w_{t^{*}}-w_{t})^{T}W_{V}(v^{*}-v_{o})>\epsilon_{1}+\epsilon_{2}$
2) When querying with a semantically equivalent key $k_{s}\in K_{s}$, by Lemma 4.1, the retrieved value is:
$v=k_{s}^{T}C^{-1}k^{*}\cdot v^{*}=\beta_{s,*}\cdot v^{*}$
where $\beta_{s,*}$ represents the whiten similarity between $k_{s}$ and $k^{*}$.
3) For robust editing, this retrieved value must maintain the prediction gap:
$(w_{t^{*}}-w_{t})^{T}W_{V}(\beta_{s,*}\cdot v^{*})>\epsilon_{1}+\epsilon_{2}$
4) Rearranging terms:
$(w_{t^{*}}-w_{t})^{T}W_{V}(k_{s}^{T}C^{-1}k^{*})\cdot v^{*T}>\epsilon_{1}+\epsilon_{2},\quad$
$\forall k_{s}\in K_{s}$
This inequality must hold for all semantically equivalent keys $k_{s}\in K_{s}$, establishing our robustness requirement.
A.4 Proof of Lemma 4.7
The specificity requirement ensures edits do not affect unrelated knowledge. We derive this as follows:
1) Consider an unrelated key $k_{o}\notin K_{s}$ with original target $t_{n}$. The corresponding output embedding is $w_{n}$.
2) To preserve specificity, the edit should not significantly alter predictions for unrelated inputs:
$(w_{n}-w_{t^{*}})^{T}W_{V}(k_{o}^{T}C^{-1}k^{*})\cdot v^{*T}<\epsilon_{3}$
3) This constraint must hold for:
  • All unrelated keys $k_{o}\notin K_{s}$
  • All possible target embeddings $w_{n}\in W$
4) Therefore, our specificity requirement is:
$(w_{n}-w_{t^{*}})^{T}W_{V}(k_{o}^{T}C^{-1}k^{*})\cdot v^{*T}<\epsilon_{3},$
$\forall k_{o}\notin K_{s},\forall w_{n}\in W$
This establishes the formal criterion for maintaining specificity in knowledge editing. The requirement ensures that edits remain localized to the intended knowledge while not affecting unrelated retrievals.
@@ B.1 Data Construction
The irrelevant long contexts are extracted from the Wikitext-103 dataset (Merity et al., 2016). The shuffled tokens are generated via sampling different word ordering. Finally, we filter the samples that are not present in the current LLM, that is, given the prefix, the target tokens are not predicted by the LLMs with the top-1 probabilities. We sample 100 samples for validation and 400 samples for test. To evaluate in-domain and out-of-domain robustness, we split the all three kinds of robustness queries in a 50-50 manner. For each sample, we have 5 in-domain queries and 5 out-of-domain queries.
B.2 Analyzing Dissimilar Keys
In Section 5.2, for each subject in CounterFact, we compute the dot product for each pair of keys of a subject’s rephrases. We utilize the inputs to the FFN’s down projection of layer 5 of LLaMA-2 as our keys, consistent with previous ROME experiments. Additionally, we include the dot product values of randomly sampled keys as a baseline for comparative analysis. We normalize the whiten similarity by the similarities between the subject itself.
B.3 Details of Training REP